I already finished this work but I would love to double-check my work. Thanks!47 Calculate the mL per hr for this weight-based infusion. A provider orders a continuous heparin infusion to begin at 15 uni52. Calculate the mL per hr flow rate range for this IV drip: A provider orders an argatroban IV drip to begin at 2 mcg per k
Answer #1

ANSWER.

47) Calculation of ml/hr.

             ml/hr = Volume(ml) / Time(hr).

     Calculation of volume(ml)

          Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

Here

      Ordered dose = 15 unit/kg/hr.

      Weight of the patient = 93.8 kg.

            Desired dose(D) = 15 * 93.8

                                        = 1407 unit.

            Dose in hand(H) = 40000 unit.

                  Quantity(Q) = 500 ml

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 1407/40000 * 500

                                                      = 0.035 * 500

                                                      = 17.5 ml.

                          Time(hr) = 1 hr.

        So

                  ml/hr = 17.5

                           = 18 ml/hr.

48) Calculation of infusion rate(ml/hr).

             ml/hr = volume(ml) / Time(hr).

        Calculation of volume.

               Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

    Here

          Ordered dose = 3 mcg/kg/min.

           Weight of the patient = 85 kg.

      So

          Desired dose(D) = 3 * 85

                                      = 255 mcg.

          Dose in hand(H) = 50 mg.

   Convert mg to mcg

                 1 mg = 1000 mcg.

        Dose in hand(H) = 50 * 1000

                                    = 50000 mcg.

                 Quantity(Q) = 250 ml

As per above formula

         Volume(ml) of drug required = D/H * Q

                                                      = 255/50000 * 250

                                                      = 0.005 * 250

                                                      = 1.3 ml.

                        Time = 1 min.

        Convert ml/min to ml/hr by multiply with 60

                              ml/hr = 1.25 * 60

                                        = 75 ml/hr.

49) Calculation of infusion rate(ml/hr).

            ml/hr = Volume(ml) / Time(hr).

      Calculation of volume(ml).

           Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

    Here

          Ordered dose = 0.015 mg/kg/min.

          Patient weight = 149 pound.

     Convert pound to kg

              1 pound = 0.454 kg.

        Patient weight = 149 * 0.454

                                = 67.6 kg.

        Desired dose(D) = 0.015 * 67.6

                                    = 1.01 mg.

        Dose in hand(H) = 200 mg.

              Quantity(Q) = 500 ml.

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 1.01 / 200 * 500

                                                     = 0.005 * 200

                                                     = 1 ml

                                  Time = 1 min.

    Convert ml/min to ml/hr by multiply with 60

                    ml/hr = 1 * 60

                             = 60 ml/hr.

50) Calculation of ml/hr.

               ml/hr = Volume(ml) / Time(hr).

         Calculation of volume.

             Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

      Here

           Desired dose(D) = 2 mg.

           Dose in hand(H) = 900 mg.

                Quantity(Q) = 500 ml.

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 2/900 * 500

                                                      = 0.002 * 500

                                                      = 1 ml.

                               Time = 1 min.

    Convert ml/min to ml/hr by multiply with 60

                ml/hr = 1 * 60

                          = 60 ml/hr.

51) Calculation of ml/hr.

             ml/hr = Volume(ml) / Time(hr).

      Calculation of volume.

           Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

   Here

      Desired dose(D) = 5.5 mg.

      Dose in hand(H) = 2000 mg (2g).

           Quantity(Q) = 500 ml

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 5.5/2000 * 500

                                                    = 0.003 * 500

                                                    = 1.5 ml

                            Time = 1min.

   Convert ml/min to ml/hr by multiply with 60

                   ml/hr = 1.5 * 60

                            = 90 ml/hr.

52) Calculation of infusion rate.

                ml/hr = Volume(ml) /Time(hr).

      Calculation of volume.

              Volume = Desired dose(D)/Dose in hand(H) * Quantity(Q)

   Here

           Ordered dose = 2 mcg/kg/min to 10 mcg/kg/min.

           Weight of the patient = 88 kg.

       Desired dose(D)for this patient = 2 * 88 to 10 * 88

                                                          = 176 mcg to 880 mcg.

              Dose in hand(H) = 250 mg.

   Convert 250 mg to mcg

              1 mg = 1000 mcg.

    So

            Dose in hand(H) = 250 * 1000

                                        = 250000 mcg.

                     Quantity(Q) = 250 ml

As per above formula

    Volume(ml) of drug required = 176/250000 * 250 to 880/250000 * 250

                                                  = 0.0007 * 250 to 0.004 * 250

                                                  = 0.18 ml to 1ml

                             Time = 1 min

     Convert ml/min to ml/hr by multiply with 60

               ml/hr = 0.18 * 60 to 1 * 60

                         = 10.8 ml to 60 ml

                         = 11 ml to 60 ml.

53) Recommended minimum and maximum dose for the patient.

               Recommended dose = 4 to 5 mcg/kg/day.

               Weight of the patient = 125 lbs.

    Convert lbs to kg for calculation

              1 lbs = 0.454 kg

       Weight of the patient = 125 * 0.454

                                         = 56.8 kg.

Minimum dose required for the patient = 4 * 56.8

                                                               = 227.2

                                                               = 227 mcg per day.

Maximum dose required for this patient = 5 * 56.8

                                                                = 284 mcg per day.

54) Calculation of ml/hr.

               ml/hr = Volume(ml) / Time(hr).

       Calculation of volume.

             Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

   Here

        Desired dose(D) = 1400 units.

        Dose in hand(H) = 30000 units.

             Quantity(Q) = 500 ml.

As per above formula

      Volume of drug required = D/H * Q

                                             = 1400/30000 * 500

                                             = 0.046 * 500

                                             = 23 ml.

                       Time(hr) = 1hr.

         So

                       ml/hr = 23 ml/hr.

55) Calculation of volume(ml) of drug required.

              Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

     Here

          Desired dose(D) = 75 mg.

          Dose in hand(H) = 50 mg.

             Quantity(Q) = 1 ml.

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                     = 75/50 * 1

                                                     = 1.5 ml.